What is 3/5 times 3/10
Found 2 solutions by solver91311, RoRo Jaxson:
Answer by solver91311(24713)



To multiply two fractions, multiply the numerator times the numerator and the denominator times the denominator.

John

Egw to Beta kai to Sigma
My calculator said it, I believe it, that settles it
The Out Campaign: Scarlet Letter of Atheism



Answer by RoRo Jaxson(21)

Ok, so first you have to make the denominators the same.
3/5=6/10
so 6/10 * 3/10
=18/100=9/50